An aqueous solution of H 2 S is added to an aqueous solution of NaOH and a reaction occurs. 3.7x10 -

4 mol of OH –1 ions are measured in the 550mL solution, calculate the number of moles of H 3 O +1 ions.

Remember, Molarity = mol/L. Calculate the concentration of the H + ions. Is the solution more acidic

or basic?


1
Expert's answer
2022-06-13T09:51:04-0400

Solution:

Molarity = Moles / Liters of solution

Moles of OH− ions = 3.7×10−4 mol

Liters of solution = (550 mL) × (1 L / 1000 mL) = 0.55 L

Therefore,

[OH−] = (3.7×10−4 mol) / (0.55 L) = 6.73×10−4 M


We can convert between [OH−] and pOH using the following equations:

pOH = −log[OH−]

pOH = −log(6.73×10−4) = 3.17

pOH = 3.17


For any aqueous solution at 25∘C:

pH + pOH = 14

pH = 14 − pOH = 14 − 3.17 = 10.83

pH = 10.83

Any solution that has a pH greater than 7 is considered basic


We can convert between pH and [H+] using the following equations:

[H+] = 10−pH

[H+] = 10−10.83 = 1.48×10−11

[H+] = 1.48×10−11 M


H+ = H3O+

[H3O+] = [H+] = 1.48×10−11 M

Moles of H3O+ ions = [H3O+] × Liters of solution

Moles of H3O+ ions = (1.48×10−11 M) × (0.55 L) = 8.14×10−12 mol

Moles of H3O+ ions = 8.14×10−12 mol


Answer:

The number of moles of H3O+ ions is 8.14×10−12 mol

The concentration of hydrogen ions (H+) is 1.48×10−11 M

The solution is basic